OpenStudy (anonymous):

A helicopter is ascending vertically with a speed of 7 m/s. At a height of 139 m above the Earth, a package is dropped from a window. How much time does it take for the package to reach the ground? and what will its speed be

OpenStudy (anonymous):

when the package dropped, it has initial velocity upward. firs of all, you should find how much time need to be for stopping. You can use Vfin=vin + g*t After that, you should find how high it goes up by using Vfin^2 = Vin^2 + 2*g*y Then, the package is in free falling. Y=Vin*t + (gt^2)/2 you should find time t from here and add it that you find t first. For final speed you can use Vfin = Vin + g*t (where t is total time t )